OpenStudy (anonymous):

h(u)=square (9-u^2) I want to get the derivative. Can't I take the 9-u^2 out of the square and make it 3-u?

OpenStudy (anonymous):

o yes it can be written as (3-u)(3+u)

OpenStudy (anonymous):

I thought so. Thanks

OpenStudy (amistre64):

no... it isnt a square

OpenStudy (amistre64):

(3-u) (3-u) does NOT equal (9-u^2) it equals (9 -6u +u^2)

OpenStudy (amistre64):

try using trig substitutions... sqrt(9 - u^2); sqrt(9 - (3sin)(3sin)) ; u=3sin sqrt(9 -9sin^2) = sqrt(9(1-sin^2)) = sqrt(9(cos^2)) = 3cos

OpenStudy (amistre64):

the derivative of 3 cos is -3sin right? plug in your sin value now..

OpenStudy (amistre64):

sin = u/3 ..thats what I meant :)

OpenStudy (amistre64):

-3(u/3) = -u if I did it right....

OpenStudy (amistre64):

or am i confusing integrals and derivatives.... one of these days Ill know when Im right :)

OpenStudy (amistre64):

I get... -1/sqrt(9-u^2) is that better?

OpenStudy (amistre64):

-u/sqrt(9-u^2) ??

OpenStudy (anonymous):

(9-u^2) is not equals (9 -6u +u^2) (9-u^2) is equal to (3-u )(3+u)

OpenStudy (amistre64):

and the sqrt((3-u)(3+u)) is not (3-u)...

OpenStudy (anonymous):

formula is a^2 - b^2 = (a+b)( a-b)

OpenStudy (amistre64):

Is the sqrt(9-u^2) = (3-u)? No.

OpenStudy (amistre64):

I might be wrong on everything else I typed, but I know im right on this one :)

OpenStudy (anonymous):

the sqrt(9-u^2) is not equal to (3-u) but sqrt(9-u^2) is equal to (3-u)(3+u)

OpenStudy (amistre64):

How so? 9-u^2 is equal to (3-u)(3+u) sqrt(9-u^2) is NOT equal to (3-u) square (3-u) and get (9-u^2) for me, just try it.... try it.... cant do it...

OpenStudy (amistre64):

is [(3-u)(3+u)]^2 equal to..... where was I?

OpenStudy (radar):

Couldn't you just square \[(9-u ^{2})^{2}\]

OpenStudy (amistre64):

maybe..... just maybe...... are we allowed to? :)

OpenStudy (anonymous):

[(3-u)(3+u)]^2 equal to (3-u)^2 (3+u)^2

OpenStudy (radar):

Wouldn't that give you \[u ^{4}-18u ^{2}+81\]

OpenStudy (radar):

Then do the derivative\[4u ^{3}-36u\]

OpenStudy (amistre64):

.... im sure wali is right about something in this....

OpenStudy (amistre64):

do we have to unsquare the derivative?

OpenStudy (anonymous):

ok u win the game but i m not wrong

OpenStudy (amistre64):

lol ... but I was gonna concede :)

OpenStudy (amistre64):

maybe

OpenStudy (amistre64):

"Can I take the 9-u^2 out of the square and make it 3-u?" What was the answer?

OpenStudy (anonymous):

no no ...

OpenStudy (amistre64):

I chose "no" as well, but im not a reliable witness :)

OpenStudy (radar):

Then simplifying getting\[1/4u[(u-3)(u+3)]\]

OpenStudy (anonymous):

ok by we will meet again because i have to prepare for tomorrow calculus paper by